# From User Preferences to Base Score Extraction Functions in Gradual Argumentation

### v1

@@ 3.5. Design Choices for Base Score Extraction Functions
The presented axioms offer the necessary conditions that the base score extraction functions must satisfy. Despite that, determining the base scores is still a flexible and context-dependent challenge. This section determines different design choices that must be considered when designing a base score extraction function.
@@ 3.5.1. Setting the Range
A choice that must be made is considering where the base scores of the most and least preferred arguments should be placed. Establishing them in 1 and 0, respectively, determines a strict and high importance of the preferences, but also allows for possible saturations from incoming supports or attacks, which would provoke information loss, an existing concern in PAFs. For a further investigation, we define the base score limits.
@@
The range is constrained within the range of the gradual semantics, i.e. $\mathbb{I}_{\sigma}\in[0,1]$. It must be satisfied that $0\leq\bot\leq\top\leq 1$. A base score extraction function will be considered normalised if its limits are set to $\top=1$ and $\bot=0$.
@@
(Base Score Normalisation) A base score extraction function $\nu$ satisfies Base Score Normalisation iff $\top=1$ and $\bot=0$.
@@
(Base Score Centralisation) A base score extraction function $\nu$ satisfies Base Score Centralisation iff $\top=1$-$\bot$.
The range influences the base scores compression. For example, if the range is too low, the base scores would be very similar. Hence, the preferences of the arguments would not be properly reflected. The following example represents the concern.
@@
Given the arguments $a,b,c,d,e,f\in\mathcal{X}$, and the preference ordering: $c\simeq f\succ b\simeq e\succ a\simeq d$. If the limits were set to $\top=0.9$ and $\bot=0.85$, the base scores could be placed only between $[0.85,0.9]$ (e.g., $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.875,\tau(a)=\tau(d)=0.85$). Alternatively, if $\bot=0.1$, the base scores could be placed in the range $[0.1,0.9]$ (e.g., $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.5,\tau(a)=\tau(d)=0.1$).
Although the base scores from the previous examples satisfy all the axioms, in the first case, they may not differentiate the preferences enough for some applications, since the difference between them is too low. As seen in the second case, decreasing the bottom limit results in a more expanded range of base scores in which the preference relations are more relatable.
@@ 3.5.2. Preference Regularity
In many real-world cases, users express their preferences in a consistent or regular way, meaning that the perceived difference in importance between two consecutively preferred arguments is roughly uniform across the ordering. For instance, if a user slightly prefers argument $a$ over $b$, and $b$ over $c$, the gap in importance between $a$ and $b$ is comparable to that between $b$ and $c$. Assuming such regularity simplifies the computation of base scores.
Before introducing the property, it is necessary to define the concept of adjacent arguments.
@@
Let $a,b\in\mathcal{X}$ be two arguments, and $\succeq_{p}$ a preference ordering over the arguments in $\mathcal{X}$, the arguments $a,b$ are adjacent in the preference ordering iff one of the following conditions holds:
@@
(Base Score Relation Regularity) A base score extraction function $\nu$ satisfies Base Score Relation Regularity iff for a given set of arguments $\mathcal{X}$, for any preference relation $p\in\{\succ\!\succ,\succ,\simeq\}$, and any two pairs of adjacent arguments in the preference ordering, $(a,b)$ and $(c,d)$, with the same preference relation, then $\nu_{p}(a)-\nu_{p}(b)=\nu_{p}(c)-\nu_{p}(d)$.
@@
(cont.) Given the previous arguments $a,b,c,d,e,f\in\mathcal{X}$, and the preference ordering: $c\simeq f\succ b\simeq e\succ a\simeq d$. If the limits were $\top=0.9$ and $\bot=0.1$, if the base score extraction function satisfies Base Score Relation Regularity, then the base scores would be $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.5,\tau(a)=\tau(d)=0.1$. Otherwise, the base scores $\tau(b)$ and $\tau(c)$ could be arbitrarily placed between 0.1 and 0.9.
If this regularity condition is not enforced, one could theoretically encode preference intensity through irregular spacing between arguments rather than using a separate much greater relation. Yet, this flexibility makes the base score extraction process more difficult, as there is no clear rule for assigning consistent numerical differences.
Another property relating to the regularity of the structure is that if new arguments are added to the framework, and these are equally preferred to existing ones, the base scores of the rest of the arguments will not change.
@@
(Base Score Preference Stability) Let $\succeq_{p}$ be a preference ordering for arguments in $\mathcal{X}$, and a new set of arguments $S$ such that $\mathcal{X}^{\prime}=\mathcal{X}\cup S$, a base score extraction function $\nu$ satisfies Base Score Preference Stability iff for the new preference ordering $\succeq_{p^{\prime}}$ such that $\forall s\in S$, $\exists a\in\mathcal{X}$ such that $a\simeq s$, then $\nu_{p}(\mathcal{X})=\nu_{p^{\prime}}(\mathcal{X}^{\prime}\setminus S)$.
@@ 3.5.3. Preferences ratios
While Preference Regularity assumes that all consecutive preferences correspond to equal differences in base scores, this may not always reflect how people actually weigh options. In practice, users often express varying intensities in their preferences. For instance, one argument may be only slightly preferred over another, while another is much more important. To model such heterogeneous gaps, we introduce the concept of preference ratios, which quantify how much stronger a much greater preference should be compared to a regular preference ($\succ$). This ratio provides a systematic way to translate qualitative intensity (e.g., “much more important”) into quantitative differences between base scores.
@@
(cont. ) Now the preference ordering becomes: $c\simeq f\succ\!\succ b\simeq e\succ a\simeq d$. From asking the human, the ratio of the much greater preference is set to 3. The base score extraction function limits are set to 0.9 and 0.1. Then, the base scores could be set to $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.3,\tau(a)=\tau(d)=0.1$. Here, the difference between preferred arguments is 0.2, while for much preferred arguments is 0.6, hence the ratio is respected.
@@
These examples highlight that the preference ratio serves as a tuning parameter controlling how sensitively the system reacts to much greater preferences. A well-chosen ratio helps capture meaningful differences in user priorities while maintaining numerical stability in base score extraction. The combination of these three different design choices thus gives a formal representation of how base score extraction functions may be personalised to different settings.
@@ 4. Concrete Base Score Extraction Functions
In this section, two base score extraction functions are introduced. Those functions are based on the preference ordering of the arguments and allow for the presented design choices. An algorithm is developed to compute the base scores. Some examples of its use are shown using the running example. By definition, as the base score extraction functions are based on preference orderings, then, scenarios with only one argument are not considered.
Since preference orderings are monotonic, considering a descending preference ordering (from the most preferred to the least preferred), the base score extraction function should also be monotonically decreasing.
@@
A base score extraction function is monotonically decreasing (increasing) iff given two arguments $a,b$ where $a\succeq b$, then $\nu_{p}(a)\leq\nu(b)$ ($\nu_{p}(a)\geq\nu(b)$, respectively).
@@
A monotonically increasing base score extraction function violates Axiom 1 for a descending preference ordering. [^4] The proof of this proposition is found in Appendix A.
Creating a monotonically decreasing function directly from the preference ordering is possible. The following method consists of assigning a distance $d$ to each argument according to the preference ordering and then normalising it between the desired range. The distance is relative to the most preferred argument.
@@
(Adaptable range) Our first base score extraction function allows for the range to be set and is defined as follows:
$\nu_{1}(\top,\bot,D,d)=\bot+(\top-\bot)\cdot\frac{D-d}{D-1}$ (1)
where $D$ is the highest distance, $d$ is the current distance, $\bot$ is the base score of the least preferred argument, and $\top$ is the base score for the most preferred argument.
The edges of this base score function correspond to when $d=1$ and $d=D$, in which the function becomes $v_{1}=\top$ and $v_{1}=\bot$ respectively. Then, with the restriction $0\leq\bot\leq\top\leq 1$, both values are ranged in $[0,1]$. Given these constraints, the function is defined inside $[0,1]$.
The next base score extraction function consists of setting the compression of the base scores and the distances towards the edges.
@@
(Adaptable squeezing and distancing) Our next base score extraction function allows a parameter to determine how compressed the base scores are to be set, and another to determine how distant the edges are, and is defined as:
$\nu_{2}(\alpha,\beta,D,d)=\frac{D-d+\alpha}{D-1+\beta}$ (2)
where $\beta\in[0,+\infty)$ is a constant related to the compression of the base scores, and $\alpha\in[0,+\infty)$ is a constant related to the distance to the edges. Intuitively, the bigger $\beta$ is, the more compressed the base scores are, while the bigger $\alpha$ is, the closer the preferred argument will be to 1.
It is necessary that $\alpha\leq\beta$, otherwise the base scores would be greater than 1. With $\beta=D-1$, the spread of the base scores is compressed in a range of 0.5. This base score extraction function is more suitable than the first one when the number of arguments is unknown or very likely to change. The first one allows for the range to be set directly, which provides more control over the edges.
Algorithm 1 shows how to extract the base scores from a set of arguments $\mathcal{X}$ and a preference ordering $\succeq_{p}$. First, the algorithm assigns $d=1$ to the most preferred argument, then, for the following arguments in the ordering, $d$ increases by $\delta$ between preferred relations and by $\Delta$ for the much preferred relationship.
@@
After assigning this distance, the base scores are obtained using one of the proposed base score extraction functions. An example applying these distances in Algorithm 1 is shown:
@@
1:A set of arguments $\mathcal{X}$, its preference ordering $\succeq_{p}$, the increase for greater preference relation $\delta$, for much greater $\Delta$, the constants $c_{1},c_{2}$ ($\top,\bot$ or $\alpha,\beta$) depending on the base score extraction function $\nu(c_{1},c_{2},D,d)$.
2:Set of argument base scores $\mathcal{T}$
@@
8:for each argument $a\in\mathcal{X}$ ordered by $\succeq_{p}$ do
9:  if $a$ is not the most preferred then
10:   $b\leftarrow$ the previous argument in $\succeq_{p}$
@@
16:     $d\leftarrow d$
17:  $\tau(a)\leftarrow\nu(c_{1},c_{2},D,d)$
18:  $\mathcal{T}\leftarrow\tau(a)$
@@
(cont.) The final strengths for the decision arguments are computed based on the previous example base scores, using the Quadratic Energy Model Potyka_kr18, which results in: $\sigma(D_{1})=0.54$ and $\sigma(D_{2})=0.40$, then the robot decides to move slowly. Note that the user prefers arguments $c$ and $f$, which increase the strength of the slow option and decrease the strength of the fast option, respectively. The arguments $b$ and $e$ increase and decrease the strength of the slow option, respectively. Finally, the arguments $a$ and $d$ aim to make the robot go fast. The algorithm’s selection of moving the robot slowly accurately adapts to the user’s preferences.
These base score extraction functions offer different benefits. First, both are flexible, allowing the base scores to be set in a desired range and proportionally adjusting the difference of base scores between arguments. Additionally, the parameters are easily modifiable, allowing corrections and learning from human feedback. Other functions could be used, e.g., an exponentially increasing or decreasing function, but there is a high chance that they do not adjust properly to a user’s preferences. For example, given five arguments $a,b,c,d,e\in\mathcal{X}$ ordered as: $a\succ b\succ c\succ d\succ e$, an exponentially decreasing function could set the base scores to $\tau(a)=1$, $\tau(b)=0.9$, $\tau(c)=0.75$, $\tau(d)=0.5$, and $\tau(e)=0.1$. Assuming that the distribution of the base scores follows that function is most probably incorrect. Allowing the user to confirm the ordering as: $a\succ b\succ c\succ d\succ\!\succ e$, the base scores could be set to $\tau(a)=1$, $\tau(b)=0.866$, $\tau(c)=0.677$, $\tau(d)=0.5$, $\tau(e)=0.1$. These base scores are similar to the previous ones, but the confirmation of the large gap between arguments $d$ and $e$ provides a more accurate representation of the real preferences.

### v3

@@ 3.5. Design Choices for Base Score Extraction Functions
The presented axioms offer the necessary conditions that the BSEFs must satisfy, but determining the base scores is still a flexible and context-dependent challenge. This section outlines different design choices that must be considered when designing a BSEF.
@@ 3.5.1. Setting the Range
A choice that must be made is considering where the base scores of the most and least preferred arguments should be placed. Establishing them as 1 and 0, respectively, determines a strict and high importance of the preferences, but also allows for possible saturations from incoming supports or attacks, which would provoke information loss, an existing concern in PAFs. For a further investigation, we define the base score limits.
@@
The range is constrained to that of the gradual semantics, i.e. $[0,1]$. Namely, it must be satisfied that $0\leq\bot\leq\top\leq 1$. A BSEF will be considered normalised if its limits are set to $\top=1$ and $\bot=0$:
@@
(Base Score Normalisation) A BSEF $\nu$ satisfies Base Score Normalisation iff $\top=1$ and $\bot=0$.
@@
(Base Score Centralisation) A BSEF $\nu$ satisfies Base Score Centralisation iff $\top=1$-$\bot$.
The range influences the base scores’ compression. For example, if the range is too low, the base scores would be very similar. Hence, the preferences of the arguments would not be properly reflected. The following example illustrates this concern.
@@
Given the arguments $a,b,c,d,e,f\in\mathcal{X}$, consider the preference ordering: $c\simeq f\succ b\simeq e\succ a\simeq d$. If the limits were set to $\top=0.9$ and $\bot=0.85$, the base scores could be placed only between $[0.85,0.9]$ (e.g., $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.875,\tau(a)=\tau(d)=0.85$). Although these base scores satisfy all the axioms, they may not sufficiently differentiate the preferences enough for some applications, since the difference between them is too low.
@@ 3.5.2. Preference Regularity
In many real-world cases, users can express their preferences regularly, meaning that the perceived difference in importance between two adjacent preferred arguments is roughly uniform across the ordering. For instance, if a user slightly prefers argument $a$ over $b$, and $b$ over $c$, the gap between $a$ and $b$ is comparable to that between $b$ and $c$. Assuming such regularity simplifies the computation of base scores. Before introducing the property, we define the notion of adjacent arguments.
@@
Let $a,b\in\mathcal{X}$ and $\succeq$ be a preference ordering. The arguments $a,b$ are adjacent in $\succeq$ iff one of the following conditions holds:
@@
(Base Score Relation Regularity) A BSEF $\nu$ satisfies Base Score Relation Regularity iff for any preference relation $\succeq$ and any two pairs $(a,b)$ and $(c,d)$ of adjacent arguments in $\succeq$, then for $\tau=\nu(\succeq)$ it holds that $\tau(a)-\tau(b)=\tau(c)-\tau(d)$.
@@
(Example 3.1 Cont.) Given the preference ordering: $c\simeq f\succ b\simeq e\succ a\simeq d$, let the limits be $\top=0.9$ and $\bot=0.1$. If the BSEF $\nu$ satisfies Base Score Relation Regularity, then any base score function obtained from $\nu$ would give base scores $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.5,\tau(a)=\tau(d)=0.1$. Otherwise, the base scores $\tau(b)$ and $\tau(c)$ could be arbitrarily placed between 0.1 and 0.9.
If this regularity condition is not enforced, preference intensity could be encoded via irregular spacing between arguments rather than using a separate much greater relation. Yet, this flexibility makes the base score extraction process more difficult, as there is no clear rule for assigning consistent numerical differences.
Another property is that adding new arguments equally preferred to existing ones does not change the base scores of the other arguments.
@@
(Base Score Preference Stability) Let $\succeq$ be a preference ordering, $S$ a set of arguments such that $S\cap\mathcal{X}=\emptyset$, $\mathcal{X}^{\prime}=\mathcal{X}\cup S$, and $\succeq^{\prime}$ be a preference ordering over $\mathcal{X}^{\prime}$ such that $\forall s\in S$, $\exists a\in\mathcal{X}$ such that $a\simeq s$ and, $\forall a,b\in\mathcal{X}$, $a\succeq^{\prime}b$ iff $a\succeq b$. Then BSEFs $\nu$ (as per Definition 3) and $\nu^{\prime}$ (mapping a set of preference orderings over $\mathcal{X}^{\prime}$ to a set of base scores for arguments in $\mathcal{X}^{\prime}$) satisfy Base Score Preference Stability iff for $\tau=\nu(\succeq)$ and $\tau^{\prime}=\nu^{\prime}(\succeq^{\prime})$, for all $x\in\mathcal{X}$, it holds that $\tau(x)=\tau^{\prime}(x)$.
@@ 3.5.3. Preferences ratios
Users often express varying intensities in their preferences. For instance, one argument may be only slightly preferred over another, while another is much more important. To model such heterogeneous gaps, we introduce the concept of preference ratios, which quantify how much stronger a much greater preference should be compared to a regular preference ($\succ$). This ratio provides a systematic way to translate qualitative intensity (e.g., “much more important”) into quantitative differences between base scores.
@@
(Example 3.5.2 Cont. ) Now the preference ordering becomes: $c\simeq f\succ\!\succ b\simeq e\succ a\simeq d$. From asking the human, the ratio of the much greater preference is set to 3. The BSEF limits are set to 0.9 and 0.1. Then, the base scores could be set to $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.3,\tau(a)=\tau(d)=0.1$. Here, the difference between preferred arguments is 0.2, while for much preferred arguments is 0.6, hence the ratio is respected.
@@
These examples highlight that the preference ratio serves as a tuning parameter controlling how sensitively the system reacts to much greater preferences. A well-chosen ratio helps capture meaningful differences in user priorities while maintaining numerical stability in base score extraction. The combination of these three different design choices thus gives a formal representation of how BSEFs may be personalised to different settings.
@@ 4. Concrete Base Score Extraction Functions
In this section, two BSEFs are introduced. Those functions are based on the preference ordering of the arguments and allow for the presented design choices. An algorithm is developed to compute the base scores. Some examples of its use are shown using the running example. By definition, as the BSEFs are based on preference orderings, then, scenarios with only one argument are not considered.
Since preference orderings are monotonic, considering a descending preference ordering (from the most preferred to the least preferred), the BSEF should also be monotonically decreasing.
@@
A BSEF $\nu$ is monotonically decreasing (increasing) iff for any base score function $\tau=\nu(\succeq)$, given two arguments $a,b$ where $a\succeq b$, then $\tau(a)\leq\tau(b)$ ($\tau(a)\geq\tau(b)$, respectively).
@@
A monotonically increasing BSEF violates Axiom 1 for a descending preference ordering. [^3] The proof of this proposition is found in Civit_arxiv26.
Creating a monotonically decreasing function directly from the preference ordering is possible. The following method consists of assigning a distance $d(x)$ to each argument $x\in\mathcal{X}$ according to the preference ordering and then normalising it between the desired range. The distance $d(x)$ is relative to the most preferred arguments, and the least preferred arguments are at the maximum absolute distance $D$ to the most preferred arguments.
@@
(Adaptable range) Our first BSEF allows for the range to be set and is defined as follows. For any ordering $\succeq$ and $x\in\mathcal{X}$, let $\top$ be the base score for the most preferred arguments (see Definition 5) and $\bot$ be the base score of the least preferred arguments (see Definition 6). Then:
$\nu_{1}(\succeq)(x)=\top+(\bot-\top)\cdot\frac{D-d(x)}{D-1}$ (1)
With the restriction $0\leq\bot\leq\top\leq 1$ and $D>1$, the base score functions obtained with this BSEF are well-defined, giving base scores inside $[0,1]$. The edge cases occur at $d(x)=1$ (for the most preferred arguments) and at $d(x)=D$ (for the least preferred arguments), yielding $v_{1}(\succeq)(x)=\top$ and $v_{1}(\succeq)(x)=\bot$ respectively.
The next BSEF consists of setting the compression of the base scores and the distances towards the edge cases.
@@
(Adaptable squeezing and distancing) Our next BSEF allows a parameter to determine how compressed the base scores are to be set, and another to determine how distant the edges are, and is defined as:
$\nu_{2}(\succeq)(x)=\frac{D-d(x)+\alpha}{D-1+\beta}$ (2)
where $\beta\in[0,+\infty)$ controls the compression of the base scores, and $\alpha\in[0,+\infty)$ determines the distance to the edges. Intuitively, larger values of $\beta$ yield more compressed base scores, and larger values of $\alpha$ place the most preferred argument closer to 1. Note that it is necessary that $\alpha\leq\beta$, otherwise the base scores would be greater than 1. With $\beta=D-1$, the spread of the base scores is compressed in a range of 0.5. This BSEF is better suited to cases where the number of arguments is unknown or likely to change, whereas the first direct control of the range and the edges allows for the range to be set directly, which provides more control over the edge cases.
Algorithm 1 shows how to extract the base scores from a set of arguments $\mathcal{X}$ and its preference ordering $\succeq$. We assume that there is at least one pair of arguments related by $\succ$ or $\succ\!\succ$ (i.e. $n+m>0$, for $n$ and $m$ at lines 2 and 3 respectively), otherwise, since no preferences over the arguments exist, the base scores cannot be dictated from them. First, the algorithm assigns a distance $d=1$ to the most preferred arguments. For each subsequent argument in the ordering, the assigned distance depends on the preference relation with its predecessor: if the arguments are equally preferred ($\simeq$), the distance remains unchanged. If the relationship between arguments is of strict preference ($\succ$), the distance is increased by $\delta$, and if the relation is of much preferred ($\succ\!\succ$), the distance is increased by $\Delta$.
@@
After assigning this distance, the base scores are obtained using one of the proposed BSEFs. An example applying these distances in Alg. 1 is shown:
@@
1:A set of arguments $\mathcal{X}$, its preference ordering $\succeq$, the increase for greater preference relation $\delta$, for much greater preference relation $\Delta$, the BSEF $\nu_{i}$ (one of $\nu_{1}$ or $\nu_{2}$).
2:Set of base scores $\mathcal{T}$ for the arguments in $\mathcal{X}$
@@
8:for each argument $a\in\mathcal{X}$ ordered by $\succeq$ do
9:  if $a$ is not most preferred then
10:   $b\leftarrow$ the previous argument in $\succeq$
@@
16:     $d(a)\leftarrow d(a)$
17:  $\tau(a)\leftarrow\nu_{i}(\succeq)(a)$ $\triangleright$ Apply Eq. 1 for $i=1$ or Eq. 2 for $i=2$
18:  $\mathcal{T}\leftarrow\mathcal{T}\cup\{\tau(a)$}
@@
(Example 4 Cont.) The final strengths for the decision arguments are computed based on the previous example base scores, using the QE Model semantics, which results in: $\sigma(D_{1})=0.54$ and $\sigma(D_{2})=0.40$. Then, the robot decides to move slowly. Note that the user prefers arguments $c$ and $f$, which increase the strength of the slow option and decrease the strength of the fast option, respectively. The arguments $b$ and $e$ increase and decrease the strength of the slow option, respectively. Finally, the arguments $a$ and $d$ aim to make the robot go fast. The algorithm’s selection of moving the robot slowly accurately adapts to the user’s preferences.
These BSEFs offer different benefits. First, both are flexible, allowing the base scores to be set in a desired range and proportionally adjusting the difference of base scores between arguments. Additionally, the parameters are easily modifiable, allowing corrections and learning from human feedback. Other functions could be used, e.g., an exponentially increasing or decreasing function, but there is a high chance that they do not adjust properly to a user’s preferences. For example, given $a,b,c,d,e\in\mathcal{X}$ ordered as: $a\succ b\succ c\succ d\succ e$, an exponentially decreasing function could set the base scores at $\tau(a)=1$, $\tau(b)=0.9$, $\tau(c)=0.75$, $\tau(d)=0.5$, and $\tau(e)=0.1$. Assuming that the distribution of the base scores follows that function is most probably incorrect. Allowing the user to confirm the ordering as: $a\succ b\succ c\succ d\succ\!\succ e$, the base scores could be set to $\tau(a)=1$, $\tau(b)=0.866$, $\tau(c)=0.677$, $\tau(d)=0.5$, $\tau(e)=0.1$. These base scores are similar to the previous ones, but the confirmation of the large gap between arguments $d$ and $e$ provides a more accurate representation of the real preferences.